OpenStudy (anonymous):

Find the number of solutions for the given equation for 0° ≤ θ ≤ 360°. sin2x - 2sinx - 3 = 0

OpenStudy (trojanpoem):

Assume sinx = x x^2 - 2x - 3 = 0 and factorize it

OpenStudy (trojanpoem):

You know how to factorize right ? you can factorize it like any other equation (sinx + 1)(sinx - 3) = 0 sinx = -1 sinx = 3 (not accepted) sinx can't be greater than 1 sinx = -1 so it's the 3rd, 4th quart

OpenStudy (trojanpoem):

so it's x = 360 - 90 = 270 and x = 180 +90 = 270 same answer so x = 270